# Probabilistic Forecasting with Coherent Aggregation

### v3

@@ 3.3 Learning Objective
Let $\theta$ be a model that resides in the class of models $\Theta$ defined by the model architecture. Here $\theta$ can be thought of a non-linear function mapping from the model feature space to the parameter set for all target horizons, we have
$(\bm{{\hat{\mu}}}_{[b][h],t},\;\bm{{\hat{\sigma}}}_{[b][h],t},\mathbf{{\hat{F}}}_{[b][k][h],t})=\theta({\mathbf{x}}^{(h)}_{[b][:t]},\;{\mathbf{x}}^{(f)}_{[b][t+1:t+N_{h}]},{\mathbf{x}}^{(s)}_{[b]}).$ (16)
Let $\hat{Y}_{i,\eta,t}(\theta)$ be the random variable parameterized by $\theta$. In some problems, multi-step coherent forecasts for multiple items are needed (e.g., in retail business, coherent regional demand forecasts are required for each product). Let $u$ be the index of such item within an index set $\{1,\cdots,N_{u}\}$ of interest, and let $\tilde{Y}_{u,i,\eta,t}(\theta)$ be the coherent forecast random variable for target $y_{u,i,t+\eta}$. Using the reparametrization trick [28], within the class of parameters $\Theta$ defined by the neural network architecture, we optimize for either CRPS
$\min_{\theta\in\Theta}\sum_{u,i,\eta,t}\mathrm{CRPS}\left(y_{u,i,t+\eta},\;\tilde{Y}_{u,i,\eta,t}(\theta)\right)=\sum_{u,i,\eta,t}\mathbb{E}\left[|y_{u,i,t+\eta},\;\tilde{Y}_{u,i,\eta,t}|\right]-\frac{1}{2}\mathbb{E}\left[|\tilde{Y}_{u,i,\eta,t}-\tilde{Y}^{{}^{\prime}}_{u,i,\eta,t}|\right].$ (17)
@@
$\min_{\theta\in\Theta}\sum_{u,t}\mathrm{ES}\left(\mathbb{P}_{u,t,[i][h]},\;\tilde{Y}_{u,t,[i][h]}(\theta)\right)=\sum_{u,t}\mathbb{E}\left[||y_{u,t,[i][h]}-\;\tilde{Y}_{u,t,[i][h]}||^{\beta}_{2}\right]-\frac{1}{2}\mathbb{E}\left[||\tilde{Y}_{u,t,[i][h]}-\tilde{Y}^{{}^{\prime}}_{u,t,[i][h]}||^{\beta}_{2}\right].$ (18)
We train the model using stochastic gradient descent (Adam [27]) with early stopping [54]. Appendix C contains the network’s optimization and hyperparameter selection details.
@@ 3.4 Discussion
Here, we discuss differences between DeepCoFactor (our method) with two end-to-end coherent probabilistic forecasting baselines in HierE2E [43] and DPMN [37].
@@
The HierE2E method of Rangapuram et al. [43] is “too general.” It consists of an augmented DeepVAR neural network model [13] that produces probabilistic forecasts for all time-series in the hierarchy. HierE2E claims to be more general than hierarchical forecasting, since it is designed to enforce any convex constraint satisfied by the forecasts; due to the constraining operation in the method, it has to revise the optimized forecasts. It does not leverage specifics of the hierarchical constraints, which are more structured than a general convex constraint. HierE2E produces forecast samples from Gaussian distributions for each time-series in the hierarchy, assuming independence; since the samples are not guaranteed to be hierarchically coherent, HierE2E couples samples by projecting them on the space of coherent probabilistic forecasts. Both the sampling operation [28] and the projection are differentiable, allowing the method to be trained end-to-end. HierE2E allows different distribution choices, although they are not explored in the initial paper, since Gaussians can be replaced by any distribution which can be sampled in a differentiable way, i.e., almost any continuous distribution [44, 12, 23]. In Rangapuram et al. [43], the projection operator ensures coherence, and correlations between bottom-levels are learned only by optimizing the neural network. In contrast, DeepCoFactor produces forecasts for bottom-level series only, while relying on common factors to encode correlations. This removes the need to forecast at all levels simultaneously, therefore reducing computational requirements if we are only interested in a subset of the aggregates.
@@
On the other hand, the DPMN baseline [37] is “too restrictive,” in particular as a Poisson Mixture can be prone to distribution mis-specification problems. It is known that when a probability model is mis-specified, optimizing log likelihood is equivalent to minimizing Kullback–Leibler (KL) divergence with respect to the true probabilistic distribution, KL divergence measures change in probability space, while optimizing CRPS is equivalent to minimizing the Cramer-von Mises criterion [16], which quantifies the distance with respect to the probability model in the sample space. The DeepCoFactor learning objective for the probabilistic model is resilient to distributional mis-specification [3]. Moreover DeepCoFactor can be optimized to adapt for other evaluation metrics of interest.
@@
Finally DPMN estimates the covariance among time series, but it does not leverage multivariate inputs when encoding the historical time series. Similar to other ARIMA based baselines, on specific hierarchical benchmark datasets such as Traffic, DPMN produces sub-optimal bottom-series forecasts. We improve the encoder for historical time series by adding a CrossSeriesMLP after the Temporal convolution encoder, which bridges the accuracy gap between HierE2E and our MQCNN based approach.
@@ 4 Empirical Evaluation
In this section, we present our main empirical results. First, we describe the empirical set up. Second, we evaluate the proposed model, and compare with state-of-the-art hierarchical forecast models. Third, we present ablation study results that further analyze the source of improvements on variants of the DeepCoFactor.
@@ 4.1 Setting
We analyze six qualitatively different public datasets: Labour, Traffic, Tourism-S, Tourism-L, Wiki, and Favorita, each requiring significant modeling flexibility due to their varied properties. The Favorita dataset is the largest dataset evaluated, which includes count and real-valued regional sales data, with over 340,000 series. The Tourism-S and Tourism-L datasets, report quarterly and monthly visitor numbers to Australian regions, respectively, and they are grouped by region and travel purpose. The Traffic dataset contains daily highway occupancy rates from San Francisco Bay Area, featuring highly correlated series with strong Granger causalities. The smallest Labour dataset tracks monthly Australian employment by status, gender, and geography, the series in this dataset are highly cointegrated. Lastly, the Wiki dataset summarizes daily online article views by country, topic, and access type. We provide more dataset details in Appendix D.
@@
Our main evaluation metric is the mean scaled CRPS from Eqn. 5 defined as the score described in Eqn. 19, divided by the sum of all target values. Let $\bm{l}^{(g)}$ be a vector of length $N_{a}+N_{b}$ consisting of binary indicators for a hierarchical level $g$, where for each $j\in[i]$, $l^{(g)}_{j}=1$ if aggregated series $j$ is included in hierarchical level $g$, and 0 otherwise. Then sCRPS for hierachical level $g$ is defined as
@@
We compare our method with the following coherent probabilistic methods: (1) DPMN-GroupBU [37], (2) HierE2E [43], (3) ARIMA-PERMBU-MinT [5], (4) ARIMA-Bootstrap-BU [40] and (5) an ARIMA. In addition, in Appendix G we compare our method with the following coherent mean methods: (1) DPMN-GroupBU, (2) ARIMA-ERM [4], (3) ARIMA-MinT [53], (4) ARIMA-BU, (5) an ARIMA and (6) Seasonal Naive. We use the implementation of statistical methods available in StatsForecast and HierarchicalForecast libraries [36, 14].
@@ 4.2 Forecasting Results
Data Level DeepCoFactor (crps) DeepCoFactor (energy) DPMN-GroupBU HierE2E * PERMBU-MinT ** Bootstrap-BU ARIMA (not coherent)
@@
As mentioned earlier, we compare the proposed model to the DPMN [37], the HierE2E [43], and two ARIMA-based reconciliation methods [53, 40]. Following previous work, we report the sCRPS at all levels of the defined hierarchies; see Table LABEL:tab:results. The ARIMA reconciliation results are generated using Olivares et al. [36], with confidence interval computed based on 10 independent runs. Results for HierE2E are generated based on three independent runs using hyperparameters tuned by Olivares et al. [36]. All metrics for DPMN are quoted from Olivares et al. [37] with identical experimental setting on all datasets.
As the Overall row in Table 3 shows, DeepCoFactor improves sCRPS upon the best alternative in five out of six datasets, with gains of on Favorita of $27.67\%$, on Tourism-S of $10.24\%$, on Tourism-L of $4.16\%$, on Wiki of $4.5\%$ and $54.40\%$ , Tourism-L and Traffic. And a degradation of $8\%$ over the best performing reconciliation on Labour, the smallest dataset. For five out of six datasets our model achieves the best or second best sCRPS accuracy across all the levels of the hierarchy, it is important to consider that aggregate levels are much smaller in sample size for which we prefer the bottom-level measurements as an indicator of the methods’ accuracy.
On Traffic, our model achieves remarkably better results than DPMN and HierE2E baselines, which we explain by the ability to model VAR relationships accurately consequence of smoothing the time series features before using them as inputs for other series. For the smallest dataset, Australian Labour, all deep learning models, including DeepCoFactor, showed a decline in accuracy compared to statistical baselines. We hypothesize this performance drop is due to the limited data available, as 57 monthly series may not be sufficient to effectively train complex models like deep learning architectures. Another explanation for the Labour degradation is the presence of strong trends and cycles, which can lead to artificially high correlations between the series, resulting in spurious relationships that may confuse the VAR modules.
We complement the main results with a mean forecast evaluation section in Appendix G. As we see Table 7 the sCRPS accuracy gains are mostly mirrored on the relative squared error metric (relSE) . We qualitatively show DeepCoFactor forecast distributions for a hierarchical structure in the Favorita dataset in Appendix E. We also include in Appendix E a probability-probability comparing the similarity of the empirical and forecast distributions, as it can be seen in Figure 7 the forecast distributions are qualitatively well calibrated.
@@ 4.3 Ablation Studies
To analyze DeepCoFactor’s source of improvements, we performed two ablation studies on the Traffic dataset, where we investigate first the effects of the learning objective, and second the effects of VAR inputs enabled by the CrossSeriesMLP. Here we report a summary and refer to details in Appendix F.
In the first study, we explore the effects of a cross-series MLP that mimics the vector autoregressive model, we compare the DeepCoFactor architecture with and without the cross series multilayer perceptron (CrossSeriesMLP) introduced in Eqn. 14. Such a module enables the network to share information of the series in the hierarchy with minimal modifications in the architecture. Table 5, and Figure 3 show the CrossSeriesMLP improved Traffic forecasting accuracy by 66% when compared to variants without it. We attribute the effectiveness of the VAR approach to the presence of Granger-causal relationships in the traffic intersections. The VAR augmented DeepCoFactor improves upon well-established univariate architectures, including LSTM [45], NBEATS [38, 34], NHITS [7], TFT [30] and FCGAGA [39] a spatio-temporal specialized architecture.
In the second ablation study on the learning objectives, we compared the CRPS-based optimization, as described in Eqn. 17, with the negative log-likelihood estimation for the Gaussian factor model introduced in Section 3.1, and other likelihood-estimated distributions. Table 6, and Figure 3 show that the CRPS-optimized factor model, improves forecasting accuracy by nearly 60% when compared to the log likelihood optimized.
Figure 3: Ablation studies on the Bay Area Traffic dataset: a) In highly correlated hierarchies, VAR inputs enabled by the DeepCoFactor to significantly improve over the univariate MQCNN’s accuracy. b) The factor model CRPS learning objective demonstrates clear advantages over classic negative log-likelihood. Full ablation studies described in Appendix F.
@@ 5 Conclusion
In this work, we present a novel multivariate factor forecasting model, integrated with the MQCNN neural network architecture, resulting in the Deep Coherent Factor Model Neural Network (DeepCoFactor) model. By optimizing CRPS as the learning objective, we achieve significant improvements in forecast accuracy over traditional negative log-likelihood objectives.
@@
Our model’s success on datasets with strong Granger causality is due to its vector autoregressive capabilities. However, in the presence of cointegrated series, vector autorregresive approaches can be ineffective. Developing multivariate time series methods that are robust to cointegration is a promising area for future research.
This study uses the reparametrization trick to enable alternative learning objectives like CRPS and the energy score for neural forecasting tasks and the use of coherent factor models to capture correlations among hierarchical series structures. While we focus on parametrizing the multivariate predictive distribution as a Gaussian multivariate factor model, the frameworks flexibility can also accommodate other distributions that support sample differentiability. This is of special interest for outlier quantiles that cannot be well approximated by Gaussian variables. Exciting future research directions include extending the reparameterization trick to handle discrete distributions, which could further enhance the accuracy of forecast distributions built on this framework. Another direction for future research is to extend the usage of the reparameterization trick from the learning objectives, into the hierarchical aggregation structure itself, provided the aggregation structure is done through differentiable transformations.
Finally we observed performance challenges of neural forecasting models when working with limited dataset sizes. Traditional tools like data augmentation and pretraining need to be adapted to handle multivariate time series and associated covariates. Extending these techniques to the hierarchical forecasting domain is a promising research direction.

### v4

@@ 3.3 Learning Objective
Let ${\bm{\theta}}$ be a model that resides in the class of models $\Theta$ defined by the model architecture. Here, $\theta$ can be thought of a non-linear function mapping from the model feature space to the parameter set for all target horizons, we have
$(\bm{{\hat{\mu}}}_{[b][h],t},\;\bm{{\hat{\sigma}}}_{[b][h],t},\mathbf{{\hat{F}}}_{[b][k][h],t})={\bm{\theta}}({\mathbf{x}}^{(h)}_{[b][:t]},\;{\mathbf{x}}^{(f)}_{[b][t+1:t+N_{h}]},{\mathbf{x}}^{(s)}_{[b]}).$ (16)
Let $\hat{Y}_{i,\eta,t}({\bm{\theta}})$ be the random variable parameterized by ${\bm{\theta}}$. In some problems, multi-step coherent forecasts for multiple items are needed (e.g., in retail business, coherent regional demand forecasts are required for each product). Let $u$ be the index of such an item within an index set $\{1,\cdots,N_{u}\}$ of interest, and let $\tilde{Y}_{u,i,\eta,t}({\bm{\theta}})$ be the coherent forecast random variable for the target $y_{u,i,t+\eta}$. As we explain in Appendix I, using the reparameterization trick [29], within the class of parameters $\Theta$ defined by the neural network architecture, we optimize for either CRPS
$\min_{{\bm{\theta}}\in\Theta}\sum_{u,i,\eta,t}\mathrm{CRPS}\left(y_{u,i,t+\eta},\;\tilde{Y}_{u,i,\eta,t}({\bm{\theta}})\right)=\sum_{u,i,\eta,t}\mathbb{E}\left[|y_{u,i,t+\eta},\;\tilde{Y}_{u,i,\eta,t}|\right]-\frac{1}{2}\mathbb{E}\left[|\tilde{Y}_{u,i,\eta,t}-\tilde{Y}^{{}^{\prime}}_{u,i,\eta,t}|\right].$ (17)
@@
$\min_{{\bm{\theta}}\in\Theta}\sum_{u,t}\mathrm{ES}\left(\mathbb{P}_{u,t,[i][h]},\;\tilde{Y}_{u,t,[i][h]}({\bm{\theta}})\right)=\sum_{u,t}\mathbb{E}\left[||y_{u,t,[i][h]}-\;\tilde{Y}_{u,t,[i][h]}||^{\beta}_{2}\right]-\frac{1}{2}\mathbb{E}\left[||\tilde{Y}_{u,t,[i][h]}-\tilde{Y}^{{}^{\prime}}_{u,t,[i][h]}||^{\beta}_{2}\right].$ (18)
We train the model using stochastic gradient descent (Adam [28]) with early stopping [56]. Appendix C contains details of the network’s optimization and hyperparameter selection.
@@ 3.4 Discussion
Here, we discuss differences between CLOVER (our method) with two coherent end-to-end probabilistic forecasting baselines in HierE2E [44] and DPMN [38].
@@
The HierE2E method of Rangapuram et al. [44] is “too general.” It consists of an augmented DeepVAR neural network model [14] that produces marginal forecasts for all hierarchical series. HierE2E claims to be more general than hierarchical forecasting, since it is designed to enforce any convex constraint satisfied by the forecasts; due to the constraining operation in the method, it has to revise the optimized forecasts. It does not leverage the specifics of the hierarchical constraints, which are more structured than a general convex constraint. HierE2E produces samples from independent Gaussian distributions for each time-series in the hierarchy; since the samples are not guaranteed to be hierarchically coherent, HierE2E couples samples by projecting them onto the space of coherent probabilistic forecasts. Both the sampling operation [29] and the projection are differentiable, allowing the method to be trained end-to-end; HierE2E restricts to optimize likelihood or CRPS. Both HierE2E and CLOVER can allow different distribution choices, since Gaussians can be replaced by any distribution that can be sampled in a differentiable way, i.e., almost any continuous distribution [45, 13, 24]. In Rangapuram et al. [44], the projection operator ensures coherence, and correlations between bottom-levels are learned only by optimizing the neural network. In contrast, CLOVER produces forecasts for bottom-level series only, while relying on common factors to encode correlations. This removes the need to forecast at all levels simultaneously, therefore, reducing computational requirements if we are only interested in a subset of the aggregates.
@@
On the other hand, the DPMN baseline [38] is “too restrictive,” in particular, as a Poisson mixture can be prone to distribution misspecification problems. It is known that when a probability model is misspecified, optimizing log likelihood is equivalent to minimizing Kullback–Leibler (KL) divergence with respect to the true probabilistic distribution, KL divergence measures change in probability space, while optimizing CRPS is equivalent to minimizing the Cramer-von Mises criterion [17], which quantifies the distance with respect to the probability model in the sample space. The CLOVER learning objective for the probabilistic model is resilient to distributional misspecification [4]. Moreover, CLOVER can be optimized to accommodate other evaluation metrics of interest.
@@
Finally, DPMN estimates the covariance among time series, but does not take advantage of the multivariate input when encoding historical time series. Similarly to other ARIMA-based baselines, on specific hierarchical benchmark datasets such as Traffic, DPMN produces suboptimal bottom-series forecasts. We improve the encoder for historical time series by adding a CrossSeriesMLP after the Temporal convolution encoder, which bridges the accuracy gap between HierE2E and our MQCNN-based approach.
@@ 4 Empirical Evaluation
In this section, we present our main empirical results. First, we describe the empirical setup. Second, we evaluate CLOVER and compare it with state-of-the-art hierarchical forecast models. Third, we present the results of the ablation study that further analyze the source of improvements in variants of CLOVER.
@@ 4.1 Setting
We analyze six qualitatively different public datasets: Labour, Traffic, Tourism-S, Tourism-L, Wiki, and Favorita, each requiring significant modeling flexibility due to their varied properties. The Favorita dataset is the largest dataset evaluated which includes count and real-valued regional sales data, with over 340,000 series. The Tourism-S and Tourism-L datasets, report quarterly and monthly visitor numbers to Australian regions, respectively, and they are grouped by region and travel purpose. The Traffic dataset contains daily highway occupancy rates from the San Francisco Bay Area, featuring highly correlated series with strong Granger causalities. The smallest Labour dataset tracks monthly Australian employment by status, gender, and geography, the series in this dataset are highly cointegrated. Lastly, the Wiki dataset summarizes daily views of online articles by country, topic, and access type. We provide more dataset details in Appendix D.
@@
Our main evaluation metric is the mean scaled CRPS from Eqn. 5 defined as the score described in Eqn. 19, divided by the sum of all target values. Let $\bm{l}^{(g)}$ be a vector of length $N_{a}+N_{b}$ consisting of binary indicators for a hierarchical level $g$, where for each $j\in[i]$, $l^{(g)}_{j}=1$ if aggregated series $j$ is included in the hierarchical level $g$, and 0 otherwise. Then sCRPS for the hierarchical level $g$ is defined as
@@
We compare our method with the following coherent probabilistic methods: (1) DPMN-GroupBU [38], (2) HierE2E [44], (3) ARIMA-PERMBU-MinT [6], (4) ARIMA-Bootstrap-BU [41], and (5) an ARIMA. In addition, in Appendix G, we compare our method with the following coherent mean methods: (1) DPMN-GroupBU, (2) ARIMA-ERM [5], (3) ARIMA-MinT [55], (4) ARIMA-BU, (5) an ARIMA, and (6) Seasonal Naive. We use the implementation of statistical methods available in the StatsForecast and HierarchicalForecast libraries [37, 15].
@@ 4.2 Forecasting Results
Data Level CLOVER (crps) CLOVER (energy) DPMN-GroupBU HierE2E * PERMBU-MinT ** Bootstrap-BU ARIMA (not coherent)
@@
As mentioned above, we compare the proposed model to the DPMN [38], the HierE2E [44], and two ARIMA-based reconciliation methods [55, 41]. Following previous work, we report the sCRPS at all levels of the defined hierarchies; see Table LABEL:tab:results. The reconciliation results for ARIMA are generated using HierarchicalForecast Olivares et al. [37], with each confidence interval calculated based on 10 independent runs. The results for HierE2E are generated based on three independent runs using hyperparameters tuned by Olivares et al. [37]. All metrics for DPMN are quoted from Olivares et al. [38] with identical experimental settings on all datasets.
As the Overall row in Table 3 shows that CLOVER improves sCRPS compared to the best alternative in five of six datasets, with gains of $27.67\%$ on Favorita, $10.24\%$ on Tourism-S, $4.16\%$ on Tourism-L, $4.5\%$ on Wiki and $54.40\%$ in Traffic. There is a degradation of $8\%$ in the reconciliation that performs the best on Labour, the smallest data set. For five out of six datasets, our model achieves the best or second-best sCRPS accuracy across all the levels of the hierarchy; it is important to consider that aggregate levels are much smaller in sample size for which we prefer the bottom-level measurements as an indicator of the methods’ accuracy.
In Traffic, our model achieves remarkably better results than the DPMN and HierE2E baselines, which we explain by the ability to model VAR relationships accurately because of the smoothing the time series features before using them as inputs for other series. For the smallest dataset, Australian Labour, all deep learning models, including CLOVER, showed a decline in accuracy compared to statistical baselines. We hypothesize that this performance drop is due to the limited data available, as 57 monthly series may not be sufficient to effectively train complex models like deep learning architectures. Another explanation for the Labour degradation is the presence of strong trends and cycles, which can lead to artificially high correlations between the series, resulting in spurious relationships that may confuse the VAR modules.
We complement the main results with a mean forecast evaluation section in Appendix G. As we see in Table 7 the sCRPS accuracy gains are mostly mirrored in the relative squared error metric (relSE) . We qualitatively show the CLOVER forecast distributions for a hierarchical structure in the Favorita dataset in Appendix E. We also include in Appendix E a probability-probability plot comparing the similarity of the empirical and forecast distributions. As can be seen in Figure 7, the forecast distributions are qualitatively well calibrated.
@@ 4.3 Ablation Studies
To analyze CLOVER’s source of improvements, we performed two ablation studies on the Traffic dataset, where we investigated first the effects of the learning objective and then the effects of VAR inputs enabled by the CrossSeriesMLP. Here we report a summary and refer to the details in Appendix F.
In the first study, we explore the effects of a cross-series MLP that mimics the vector autoregressive model. We compare the CLOVER architecture with and without the cross-series multi layer perceptron (CrossSeriesMLP) introduced in Eqn. 14. Such a module enables the network to share information from the series in the hierarchy with minimal modifications to the architecture. Table 5 and Figure 3 show that CrossSeriesMLP improved Traffic forecast accuracy by 66% compared to variants without it. We attribute the effectiveness of the VAR approach to the presence of Granger-causal relationships in traffic intersections. The VAR-augmented CLOVER improves upon well-established univariate architectures, including LSTM [46], NBEATS [39, 35], NHITS [8], TFT [31], and FCGAGA [40], a spatio-temporal specialized architecture.
In the second ablation study on learning objectives, we compared CRPS-based optimization, as described in Eqn. 17, with the negative log-likelihood estimation for the Gaussian factor model introduced in Section 3.1, and other likelihood-estimated distributions. Table 6 and Figure 3 show that the CRPS-optimized factor model improves forecast accuracy by nearly 60% when compared to the log-likelihood optimized model.
Figure 3: Ablation studies on the Bay Area Traffic dataset: a) In highly correlated hierarchies, VAR inputs enabled by the CLOVER to significantly improve over the univariate MQCNN’s accuracy. b) The factor model CRPS learning objective demonstrates clear advantages over classic negative log-likelihood. Full ablation studies described in Appendix F.
@@ 5 Conclusion
In this work, we present a novel multivariate factor forecasting model, integrated with the MQCNN neural network architecture, resulting in the Coherent Learning Objective Reparameterization Neural Network (CLOVER) model. Using CRPS as the learning objective, we achieve significant improvements in forecast accuracy over traditional negative log-likelihood objectives.
@@
Our model’s success on datasets with strong Granger causality is due to its vector autoregressive capabilities. However, in the presence of cointegrated series, vector autoregresive approaches can be ineffective. Developing multivariate time series methods that are robust to cointegration is a promising area for future research.
This study uses the reparameterization trick to enable alternative learning objectives such as CRPS and the energy score for neural forecasting tasks and the use of coherent factor models to capture correlations among hierarchical series structures. While we focus on parameterizing the multivariate predictive distribution as a Gaussian multivariate factor model, the framework’s flexibility can also accommodate other distributions that support sample differentiability. This is of special interest for outlier quantiles that cannot be well approximated by Gaussian variables. Exciting future research directions include extending the reparameterization trick to handle discrete distributions, which could further enhance the accuracy of forecast distributions built on this framework. Another promising direction for future research is to extend the usage of the reparameterization trick from the learning objectives, into the hierarchical aggregation structure itself, provided the aggregation structure is done through differentiable transformations.
Finally, we observed performance challenges of neural forecasting models when working with limited dataset sizes. Traditional tools like data augmentation and pretraining need to be adapted to handle multivariate time series and associated covariates. Extending these techniques to the hierarchical forecasting domain is a promising research direction.